answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

Is continuous distribution convert in discrete distribution?

If the question is asking if a continuous distribution can be converted to a discrete distribution, the answer is yes.

Your age has a continuous distribution but in most cases, the information is recorded and analysed as if it were the whole number of years - a discrete distribution.

How do you create five scores with a mean of 10 and a standard deviation of 0?

Since the standard deviation is zero, the scores are all the same. And, since their mean is 10, they must all be 10.

What are the assumptions of standard deviation?

The standard deviation is the standard deviation! Its calculation requires no assumption.

What is normal distribution in statstics?

The normal distribution is a statistical distribution.

Many naturally occurring variables follow the normal distribution: examples are peoples' height, weights.

The sum of independent, identically distributed variables - whatever their own underlying distribution - will tend towards the normal distribution as the number in the sum increases. This means that the mean of repeated measures of ANY variable will approach the normal distribution.

Furthermore, some distributions that are not normal to start with, can be converted to normality through simple transformations of the variable.

These characteristics make the normal distribution very important in statistics. See attached link for more.

When two fair sided dice are rolled what is the probability of rolling doubles or the sum of the two dice equal 6?

The sample space has 36 elements (the total number of outcomes by rolling the two dice).

There are 6 double outcomes: (1,1), (2,2), (3,3), (4,4), (5,5) and (6,6).

There are 5 outcomes whose sum is 6: (1,5), (5,1), (2,4), (4, 2), and (3,3).

The probability of rolling doubles OR the sum of 6 is 6/36 + 5/36 = 11/36.

What is probability and how does it help explain the result of crosses?

Probability is a number that describes how likely it is that an event will occur.

What are the different ways that you can arrange your windows on your desktop?

The answer will be an incredibly huge number. There are a different number of windows, each window can refer to a different application, they can be of different sizes, they can be located at different positions.

What is the theoretical probability of rolling sum of 7?

-- There are (6 x 6) = 36 possible rolls for a fair pair of 6-sided dice.

-- There are 6 ways to roll a sum of 7 :

  • 1 ... 6
  • 2 ... 5
  • 3 ... 4
  • 4 ... 3
  • 5 ... 2
  • 6 ... 1

-- So the probability is 6/36 = 1/6 = 162/3 % .

-- The probability is the probability, not the 'theoretical' probability.

What is the probablity of drawing a red card or a queen?

If you draw more than 24 cards from a standard pack, without replacement, the probability is 1. That is, it is a certainty.

The probability of the outcome for a single, randomly drawn card from a standard pack, is 7/13.

What is the probability of choosing blue marble followed by a green marble from a bag containing 5 blue 5 green and 5 red marbles?

The probability of choosing a blue marble is 5 in 15 or 1 in 3. The probability of then choosing a green marble is 5 in 14. (One is missing) Multiply the two probabilities and you get 5 in 42.(P = 0.1190... about 12%).

How many 5 card poker hands consisting of 2 kings and 2 jacks are possilbe with a 52 card deck?

There are 1584 hands possible with 2 Kings, 2 Jacks and 1 other card which is not a King or a Jack.

What is the The approximate probability of a value occurring that is greater than one standard deviation?

The answer depends on the underlying distribution. For example, if you have a random variable X, with a symmetric distribution with mean = 20 and sd = 1, then

prob(X > 1) = 1, to at least 10 decimal places.

What is the theoretical probability of the event when rolling a 12 sided dice?

The answer depends on what the "event" is. And since you have not bothered to share that crucial bit of information, I cannot provide a more useful answer.

Trending Questions
Which is farther west eastern standard centeral standard mountain standard and pacific standard? What is Brians probable diagnosis? When tossing 4 coins simultaneously find the probability that exactly 2 heads is showing? There are six pairs of black gloves and six pairs of brown gloves in a drawer In complete darkness how many gloves must you take from the drawer in order to be sure to get a pair that match? What two ports share the same estuaries as Felixstowe? What is the probability that three coin flips will have at least one head? What does the standard deviation tell us about the mean? What is the probability of two dice that have a 1 side with a one 2 sides with two on it and 3 sides with 3 on it. and not getting a 2 when you role both? The likelihood that an event will occur like tossing a coin? What is probability of throwing 9 heads in a row? A mother has one LL form color lameness and one allele for normal vision what is the probability that her Jean mitt will have the LL for color? What is the probability that the black parent of a black offspring was homozygous for black coat? A set of 1000 values has a normal distribution the mean of the data is 120 and the standard deviation is 20 how many values are within one standard deviaiton from the mean? What is the probability of dying from jumping off a building? If a bag contains 4 black 6 green and 2 yellow marbles and one marble is randomly drawn what is the probability that it is not black? How many distinguishable permutations can be made from the letters in the word gallipolis? What is the probability that 4 persons can sit in 4 different chairs? You roll two dice What is the probability that the sum of the dice is odd or the sum of the dice is greater than 6? How many 2s in deck of 52? What is the class boundary of a class ranging from 115 to 215?